# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 4.3 Longitudinal mode for a neutral fluid
Let us first consider the dispersion relation for the longitudinal mode for a neutral fluid. This argument is included in the analysis for a general U(1) charged fluid in the next subsection. However, the expressions are simpler than the charged case, and it may be useful to compare the present analysis with the previous results. ~$\tilde{\mathcal{E}}^{\rm L}=0$ is recast into a well-known form Hiscock:1985zz
@@
$+\frac{1}{3}k^{4}(3\zeta+4\eta)\kappa_{1}-\beta k^{2}P_{\beta}(\rho+P).$ (223)
@@
${\rm Im}~{}\omega\propto\kappa_{2}^{-1}\geq 0.$ (224)
While the mode is stable for the exact $\kappa_{2}=0$, the instability gets worse as $\kappa_{2}\to 0$. It was said that the $\kappa_{2}=0$ is a “peculiar singular limit” Hiscock:1985zz . This is known as hydrodynamical instability. However, such a pathological behavior is simply because of the inconsistent treatment of the gradient expansion, and Eq. (224) is outside the allowed parameter space in the effective theory. In fact, $\mathcal{O}(\varepsilon^{n>1})$ in equation of motion is understood as a higher-order derivative correction; therefore, we must treat them perturbatively Abbasi:2017tea . To be more specific, when solving the constraint equation
$\bar{\pi}^{*}_{0}=\bar{\pi}_{i}\frac{\mathcal{E}^{i}{}_{0}}{\mathcal{E}^{0}{}_{0}},$ (225)
to find Eq. (223), we need to be more careful. The dispersion relation is found by simplifying
$k_{i}k^{j}\left(\mathcal{E}^{i}{}_{j}+\frac{\mathcal{E}^{i}{}_{0}\mathcal{E}^{0}{}_{j}}{\mathcal{E}^{0}{}_{0}}\right)=0.$ (226)
At the level of EoM, one may multiply $\mathcal{E}^{0}{}_{0}$ to eliminate the denominator and find Eq. (223). However, we cannot justify the prescription at the action level. Instead, we rewrite the denominator as
$\frac{1}{2\mathcal{E}^{0}{}_{0}}=\frac{1}{-\beta\rho_{\beta}\omega^{2}+i\omega^{3}\epsilon+ik^{2}\omega\kappa}=\frac{1}{-\beta\rho_{\beta}\omega^{2}}\left(\frac{1}{1-\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}}\right).$ (227)
As we mentioned below Eq. (22), the gradient expansion is the expansion in $\varepsilon=(\rho+P)^{-1}v\nabla$ with the viscosity coefficients $v=\epsilon,\lambda,\cdots$. We have to truncate the expansion at the first order in $\mathcal{O}(\varepsilon)$:
$\frac{1}{1-\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}}=1+\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}+\mathcal{O}(\varepsilon^{2}).$ (228)
Thus, while we truncated the derivative expansion at first order in the constitutive relation, the higher order derivatives enter the dispersion relation for the NG modes when solving the constraint equations. However, the mode analysis is reliable only up to $\mathcal{O}(\varepsilon^{1})$ since the constitutive relation is first order. Hence, we rather find
@@
$-\beta k^{2}P_{\beta}(\rho+P)+\mathcal{O}(\varepsilon^{2}).$ (229)
In the first-order hydrodynamical theory, there are two energy scales: $M_{\rm SSB}$ and $M_{\rm diss.}$. $M_{\rm SSB}$ is the energy of spontaneous symmetry breaking, typically $\rho\sim M_{\rm SSB}^{4}$. For a conformal fluid, $M_{\rm SSB}\sim\beta^{-1}$. The perturbative description is not applicable for $\omega>M_{\rm SSB}$, which is typically the energy scale of a thermal fast mode. In fact, one can canonically normalize the NG mode as $\pi_{\rm c}\sim\sqrt{\rho+P}\pi$, and the action is expanded by $(\partial\pi_{\rm c}/M^{2}_{\rm SSB})^{n}$. Hence, the perturbative expansion is justified for $\omega\ll M_{\rm SSB}$. $\eta\sim M_{\rm diss.}^{3}$ is the energy scale of dissipation determined by the mean-free-path of the fast modes. The order counting parameter of the gradient expansion is $\varepsilon\sim\omega M_{\rm diss.}^{3}/M_{\rm SSB}^{4}$, and the gradient expansion is valid in the regime where $\varepsilon\ll 1$. In fact, there are several situations to realize $\varepsilon\ll 1$. The first case is a strong coupling regime where $\omega\ll E_{\rm diss.}\leq E_{\rm SSB}$: the mean-free-path is extremely small. This is a case for extremely “smooth” medium, such as quark-gluon plasma in nuclear physics. In this case, we only need to expand the solution in the following way in Eq. (229):
$\omega=\vartheta(\omega_{0}+\vartheta\omega_{1}+\mathcal{O}(\vartheta^{2})),~{}k=\vartheta k,$ (230)
where $\vartheta$ is the order counting parameter of $\omega$ and $k$. On the other hand, $\varepsilon\ll 1$ is also realized in the weak coupling regime $E_{\rm diss.}\ll\omega\ll E_{\rm SSB}$, which corresponds to
$\omega=\omega_{0}+\vartheta\omega_{1}+\mathcal{O}(\vartheta^{2}),~{}\eta=\vartheta\eta,~{}\zeta=\vartheta\zeta,\cdots$ (231)
where $\cdots$ imply the rest of 1st order coefficients, and $\vartheta$ is now the order counting parameter of $E_{\rm diss.}$. This situation is typically realized in photon-baryon plasma in the cosmic microwave background before the last scattering, where fluid is realized by slow processes such as Thomson scattering between photons and electrons.
For both scenarios, the order counting in $\vartheta$ is equivalent to $\varepsilon$, and thus, the mode analysis is the same. $E_{\rm diss.}\sim\omega\ll E_{\rm SSB}$ is included in the case of Eq. (231) with $\vartheta\to\varepsilon$. Thus, the gradient expansion describes strong and weak coupling regimes in the same way, and we do not need to distinguish these scenarios in the mode analysis. The lowest-order solution is immediately found as
$\omega_{0}=\pm c_{s}k,$ (232)
@@
$c_{s}^{2}=\frac{\partial P}{\partial\rho}=\frac{P_{\beta}}{\rho_{\beta}}.$ (233)
@@
$c_{s}^{2}$ is determined by zeroth-order hydrodynamics, i.e., thermodynamics, which is undefined at this order. The dispersion relation at first order is solved as
@@
$\omega_{1}=-\frac{2ik^{2}\eta}{3(\rho+P)}-\frac{ik^{2}}{2(\rho+P)}\left(\zeta-\frac{\partial P}{\partial\rho}\lambda_{1}\right)-\frac{ik^{2}}{2\beta P_{\beta}}\frac{\partial P}{\partial\rho}\left(\lambda_{2}-\frac{\partial P}{\partial\rho}\epsilon\right).$ (234)
@@
Thus, $\kappa_{i}$ disappears from the first-order analysis. This is reasonable as $\kappa_{i}$ is redundant freedom for a general neutral fluid: one can eliminate $\kappa_{i}$ by a field redefinition in the absence of a number flux. Note that we used Eq. (232), which corresponds to using the zeroth order hydrodynamics, just like using the on-shell solutions. However, there is no ambiguity in the use of Eq. (232). The hydrodynamical frame invariants $\tilde{\zeta}$ and $\tilde{\lambda}$ write $\omega_{1}$, so the mode analysis does not depend on the hydrodynamical frame choice. Imposing the KMS condition, one can write Eq. (234) as
@@
$\omega_{1}=-\frac{2ik^{2}\eta}{3(\rho+P)}-\frac{ik^{2}}{2(\rho+P)}\left(\zeta-2\frac{\partial P}{\partial\rho}\lambda+\left(\frac{\partial P}{\partial\rho}\right)^{2}\epsilon\right).$ (235)
@@
Now the unitarity constraint (73) implies that ${\rm Im}~{}\omega_{1}\leq 0$ if $\rho+P>0$ is satisfied. The sound speed is not affected by the first order coefficients as ${\rm Re}~{}\omega_{1}=0$. We extend this result to a general U(1) charged fluid shortly.
@@
We did not observe gapped excitations as in the case of other mode analyses in open systems Minami:2015uzo ; Hongo:2018ant . This is because the hierarchy in $\varepsilon$ dominates over the dispersion relation: mixing of $\mathcal{O}(\varepsilon^{0})$ and $\mathcal{O}(\varepsilon^{1})$ is not allowed. If we consider the “nonrelativistic” regime $k\ll\omega$, paying the price for general covariance, there may appear more richer structure. For the non relativistic strong coupling regime, $k\ll\omega\ll E_{\rm diss.}\leq E_{\rm SSB}$, let us put forward the following solution:
$\omega=\vartheta(\omega_{0}+\mathcal{O}(\vartheta^{1})),~{}k=\vartheta^{2}k.$ (236)
Then we find $\omega_{0}=0$. Note that $\mathcal{O}(\vartheta^{1})$ correction to $\omega_{0}=0$ is nothing but the relativistic regime and one finds $\omega_{1}=\pm c_{s}k$. Let us consider a slightly different case, $k\ll\omega\sim E_{\rm diss.}\leq E_{\rm SSB}$. If we naively solve the dispersion relation (229) with
$\omega=\omega_{0}+\vartheta\omega_{1}+\mathcal{O}(\vartheta^{2}),~{}k=\vartheta k,$ (237)
@@
one finds a hydrodynamical frame-dependent gapped mode, $\omega_{0}=\mathcal{O}(\varepsilon^{-1})$. Mathematically, this solution is found when the two terms in the first line in Eq. (229) balance, which is the invalid regime in the gradient expansion. The fact that a gap does not appear implies that the NG modes are massless and that dissipation is always higher order in the derivative expansion. Thus, the effective action of the hydrodynamical NG modes is subject to extra constraints compared to that of a general QFT.
@@
While $k\ll\omega$ is a nonrelativistic regime for sound waves, $P\to 0$ limit corresponds to the nonrelativistic regime of the UV sector. In this case, with Eq. (231), we find $\omega_{0}=0$, and then
$\omega_{1}=-\frac{ik^{2}(3\rho\lambda_{2}+(3\zeta+4\eta)\rho_{\beta})}{3\rho\beta\rho_{\beta}},$ (238)
@@
which can be also found when taking $P\to 0$ and $P_{\beta}\to 0$ in Eq. (234). Thus, for a vanishing pressure limit, the NG modes are purely diffusive. $\omega_{1}$ depends on $\lambda_{2}$ and $\zeta$ in this limit, which are now hydrodynamical invariants as $\tilde{\lambda}\to\lambda_{2}$ and $\tilde{\zeta}\to\zeta$ in $P\to 0$ limit. $\lambda_{1}$ and $\epsilon$ are eliminated by field redefinition, and the nontrivial first-order correction in the energy-momentum tensor is only spatial.
@@ 4.4 Longitudinal mode for a general charged fluid
For a general charged fluid, the mode analysis is more complicated than the neutral case. The dispersion relation is now 6th order algebraic equation (for a conformal U(1) charged fluid, see Ref. Taghinavaz:2020axp )
$0=\sum_{n=0}^{6}d_{n}\omega^{n},$ (239)
@@
$\left.+(\beta P_{\beta}-\mu P_{\mu})\left(3\beta n\tau_{3}-3\beta^{2}\chi_{1}(\rho+P)\right)\right),$ (240)
@@
$\left.-i(3\zeta+4\eta)\left(\beta\left(\beta^{2}(-\rho_{\beta})\chi_{1}+\beta\mu\rho_{\mu}\chi_{1}+\beta n_{\beta}\tau_{3}-\rho_{\mu}\tau_{1}\right)+n_{\mu}(\kappa_{1}-\beta\mu\tau_{3})\right)\right)$ (241)
@@
$+3(\rho+P)(\beta(\beta\chi_{1}(\beta\rho_{\beta}-\mu\rho_{\mu})+\beta\nu_{5}n_{\beta}+\rho_{\mu}\tau_{1})+n_{\mu}(\lambda_{2}-\beta\mu\nu_{5})))$ (242)
@@
$+\beta\chi_{1}(\beta\kappa_{2}\rho_{\beta}-\kappa_{2}\mu\rho_{\mu}+\rho(-\epsilon))+\beta\mathcal{P}(\lambda_{2}\nu_{3}-\nu_{1}\nu_{5}+\nu_{4}\tau_{1}-\chi_{1}\epsilon))))$ (243)
@@
$+\kappa_{1}\nu_{4}(\nu_{2}+\tau_{2})+\lambda_{1}\nu_{1}\tau_{3}+\kappa_{2}(-\lambda_{2}\nu_{3}+\nu_{1}\nu_{5}-\nu_{4}\tau_{1}+\chi_{1}\epsilon)-\nu_{2}\tau_{3}\epsilon-\tau_{2}\tau_{3}\epsilon))$ (244)
@@
$+\beta(\beta\nu_{3}(\mu\rho_{\mu}-\beta\rho_{\beta})-\nu_{1}\rho_{\mu}+\beta\nu_{4}n_{\beta}))$ (245)
$d_{6}$ $=3\beta^{2}\kappa_{2}(\nu_{3}\epsilon-\nu_{1}\nu_{4})$ (246)
As in the case of the neutral fluid, the above dispersion relation is reliable up to $\mathcal{O}(\varepsilon)$, and we need to solve the dispersion relation order by order in the gradient. In the low energy limit, regardless of strong or weak couplings, the zeroth order solution is found as
$\omega_{0}=\pm c_{s/n}k,~{}c_{s/n}^{2}\equiv\left(\frac{\partial P}{\partial\rho}\right)_{s/n}.$ (247)
The adiabatic sound velocity $c_{s/n}$ contains Eq. (233) in the neutral limit. The explicit form of $c_{s/n}$ is found in Eq. (283). Zeroth order thermodynamics is stable if $c_{s/n}^{2}\geq 0$. The first-order solution is found as
$\omega_{1}$ $=-i\Gamma_{\rm L}k^{2},$ (248)
$\Gamma_{\rm L}$ $\equiv\frac{1}{2(\rho+P)}\left[\frac{4}{3}\eta+{\bm{\zeta}}+f^{2}\left(\frac{\partial P}{\partial\rho}\right)_{s/n}{\bm{\chi}}_{1}+\beta\left(\frac{\partial P}{\partial n}\right)_{\rho}{\bm{\tau}}\right],$ (249)
where ${\bm{\zeta}}$, ${\bm{\chi}}_{1}$ and ${\bm{\tau}}$ are defined in Sec. 2.5, and we defined
$f^{2}$ $=\frac{\beta^{2}(\rho+P)^{2}((\rho+P-\mu n)(\beta n_{\beta}-\mu n_{\mu})-\beta n\rho_{\beta})^{2}}{\left(n_{\mu}(\rho+P-\mu n)^{2}-\beta n(n_{\beta}(\mu n-2(\rho+P))+n\rho_{\beta})\right)^{2}}.$ (250)
Although Eq. (249) is written in a simple form, concrete expressions for $\left(\frac{\partial P}{\partial\rho}\right)_{s/n}$ and $\left(\frac{\partial P}{\partial n}\right)_{\rho}$ are tedious, as shown in App. A.3. Note that $f^{2}\geq 0$ since the thermodynamic variables are real. ${\bm{\chi}}_{2}$ drops from the first order solution. As the bold faces are hydrodynamical frame invariants, the mode analysis does not depend on the hydrodynamical frame choice. We have substituted $\omega_{0}$ into the first-order dispersion relation to find $\omega_{1}$, which corresponds to solving the on-shell hydrodynamics for NG modes. This is why the boldfaces write the first-order dispersion relation. Importantly, there is no ambiguity in the use of the on-shell condition, in contrast to the case in Sec. 2.5. As discussed in Sec. 2.5, the KMS condition and unitarity of the UV system implies ${\bm{\chi}}\geq 0$, ${\bm{\tau}}=0$ and ${\bm{\zeta}}\geq 0$. Therefore, $\Gamma_{\rm L}\geq 0$ and stable if $\rho+P>0$. Note that we have not taken any specific hydrodynamical frame for writing down the effective action for the NG modes, and hence, this mode analysis applies to the Eckart frame as well.
@@
~$\tilde{\mathcal{E}}^{\rm L}=k^{3}\left(\frac{4}{3}\eta+\zeta+\lambda_{1}+\frac{n(\lambda_{2}+\epsilon)}{\beta n_{\beta}-\mu n_{\mu}}\right)+\mathcal{O}(\varepsilon^{2}).$ (251)

### v3

@@ 5.3 Longitudinal mode for a neutral fluid
Let us first consider the dispersion relation for the longitudinal mode for a neutral fluid. This argument is included in the analysis for a general U(1) charged fluid in the next subsection. However, the expressions are simpler than the charged case, and it may be useful to compare the present analysis with the previous results. ~$\tilde{\mathcal{E}}^{\rm L}=0$ is recast into a well-known form
@@
$+\frac{1}{3}k^{4}(3\zeta+4\eta)\kappa_{1}-\beta k^{2}P_{\beta}(\rho+P).$ (5.44)
Eq. (5.44) has been discussed in various seminal works in the past. As our approach significantly differs from others, checking the consistency of this result can be a benchmark of our detailed analysis. In fact, the same equation is found in Eq. (4.8) in a seminal work [23]. Although it was derived in a different contexts, we exactly reproduced the same result by a careful matching of conventions and use of thermodynamic relations  [^4] Readers can check the consistency with the help of the footnote 8 in Ref. [23], and our translation table 1. .
@@
${\rm Im}~\omega\propto\kappa_{2}^{-1}\geq 0.$ (5.45)
While the mode is stable for the exact $\kappa_{2}=0$, the instability gets worse as $\kappa_{2}\to 0$. It was said that the $\kappa_{2}=0$ is a “peculiar singular limit” [11]. This is known as hydrodynamical instability. However, such a pathological behavior is simply because of the inconsistent treatment of the gradient expansion, and Eq. (5.45) is outside the allowed parameter space in the effective theory. In fact, $\mathcal{O}(\varepsilon^{n>1})$ in equation of motion is understood as a higher-order derivative correction; therefore, we must treat them perturbatively [28]. To be more specific, when solving the constraint equation
$\bar{\pi}^{*}_{0}=\bar{\pi}_{i}\frac{\mathcal{E}^{i}{}_{0}}{\mathcal{E}^{0}{}_{0}},$ (5.46)
to find Eq. (5.44), we need to be more careful. The dispersion relation is found by simplifying
$k_{i}k^{j}\left(\mathcal{E}^{i}{}_{j}+\frac{\mathcal{E}^{i}{}_{0}\mathcal{E}^{0}{}_{j}}{\mathcal{E}^{0}{}_{0}}\right)=0.$ (5.47)
At the level of EoM, one may multiply $\mathcal{E}^{0}{}_{0}$ to eliminate the denominator and find Eq. (5.44). However, we cannot justify the prescription at the action level. Instead, we rewrite the denominator as
$\frac{1}{2\mathcal{E}^{0}{}_{0}}=\frac{1}{-\beta\rho_{\beta}\omega^{2}+i\omega^{3}\epsilon+ik^{2}\omega\kappa}=\frac{1}{-\beta\rho_{\beta}\omega^{2}}\left(\frac{1}{1-\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}}\right).$ (5.48)
As we mentioned below Eq. (2.22), the gradient expansion is the expansion in $\varepsilon=(\rho+P)^{-1}v\nabla$ with the viscosity coefficients $v=\epsilon,\lambda,\cdots$. We have to truncate the expansion at the first order in $\mathcal{O}(\varepsilon)$:
$\frac{1}{1-\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}}=1+\frac{i\omega^{3}\epsilon+ik^{2}\omega\kappa}{\beta\rho_{\beta}\omega^{2}}+\mathcal{O}(\varepsilon^{2}).$ (5.49)
Thus, while we truncated the derivative expansion at first order in the constitutive relation, the higher order derivatives enter the dispersion relation for the NG modes when solving the constraint equations. In fact, in Landau frame where $\kappa=\epsilon=0$, we can safely solve the constraint equation without issue, and the derivative expansion scheme in the dispersion relation is identical to that of the constitutive relation. However, the mode analysis is reliable only up to $\mathcal{O}(\varepsilon^{1})$ in a general hydrodynamical frame. We must further expand the dispersion relation to the same order of the constitutive relation. Hence, we rather find
@@
$-\beta k^{2}P_{\beta}(\rho+P)+\mathcal{O}(\varepsilon^{2}).$ (5.50)
The difference of Eq. (5.44) and (5.50) is unreliable correction in $\varepsilon$ as we only expand the constitutive relation to first order, and thus we must drop for a consistent analysis in effective theory point of view.
Next let us consider the mode analysis of Eq. (5.50). Naive mistake is to take the zero of this third order algebraic equation. One finds three roots in this analysis. One of the three modes is in fact unphysical, since it is outside the convergence radius of the derivative expansion. The truncation of this solution is not ad-hoc, as reliability of Eq. (5.50) is not valid. The consistent mode analysis should be given by expanding the solution to the series in the expansion parameter $\varepsilon$:
$\omega=\sum_{n=0}^{\infty}\varepsilon^{n}\omega_{n}.$ (5.51)
Below we determine $\omega_{n}$ order by order in $\varepsilon$. The lowest-order solution is immediately found as
$\omega_{0}=\pm c_{s}k,$ (5.52)
@@
$c_{s}^{2}=\frac{\partial P}{\partial\rho}=\frac{P_{\beta}}{\rho_{\beta}}.$ (5.53)
@@
$c_{s}^{2}$ is determined by zeroth-order hydrodynamics, i.e., thermodynamics, which is undefined at this order. This mode result implies the obvious fact that we have the free propagating sounds if the first order coefficients are small. The dispersion relation at first order is solved as
@@
$\omega_{1}=-\frac{2ik^{2}\eta}{3(\rho+P)}-\frac{ik^{2}}{2(\rho+P)}\left(\zeta-\frac{\partial P}{\partial\rho}\lambda_{1}\right)-\frac{ik^{2}}{2\beta P_{\beta}}\frac{\partial P}{\partial\rho}\left(\lambda_{2}-\frac{\partial P}{\partial\rho}\epsilon\right).$ (5.54)
@@
Thus, $\kappa_{i}$ disappears from the first-order analysis. This is reasonable as $\kappa_{i}$ is redundant freedom for a general neutral fluid: one can eliminate $\kappa_{i}$ by a field redefinition in the absence of a number flux. Note that we used Eq. (5.52), which corresponds to using the zeroth order hydrodynamics, just like using the on-shell solutions. However, there is no ambiguity in the use of Eq. (5.52). The hydrodynamical frame invariants $\tilde{\zeta}$ and $\tilde{\lambda}$ write $\omega_{1}$, so the mode analysis does not depend on the hydrodynamical frame choice. Imposing the KMS condition, one can write Eq. (5.54) as
@@
$\omega_{1}=-\frac{2ik^{2}\eta}{3(\rho+P)}-\frac{ik^{2}}{2(\rho+P)}\left(\zeta-2\frac{\partial P}{\partial\rho}\lambda+\left(\frac{\partial P}{\partial\rho}\right)^{2}\epsilon\right).$ (5.55)
@@
Now the unitarity constraint (2.70) implies that ${\rm Im}~\omega_{1}\leq 0$ if $\rho+P>0$ is satisfied. The sound speed is not affected by the first order coefficients as ${\rm Re}~\omega_{1}=0$. We extend this result to a general U(1) charged fluid shortly.
@@
We did not observe gapped excitations as in the case of other mode analyses in open systems [29, 30]. This is because the hierarchy in $\varepsilon$ dominates over the dispersion relation: mixing of $\mathcal{O}(\varepsilon^{0})$ and $\mathcal{O}(\varepsilon^{1})$ is not allowed. Gapped mode might be found in the regime, $k\ll\omega$. If we naively solve the dispersion relation (5.50) with
$\omega=\omega_{0}+\vartheta\omega_{1}+\mathcal{O}(\vartheta^{2}),~k=\vartheta k,$ (5.56)
@@
one finds a hydrodynamical frame-dependent gapped mode, $\omega_{0}=\mathcal{O}(\varepsilon^{-1})$. Mathematically, this solution is found when the two terms in the first line in Eq. (5.50) balance, which is the invalid regime in the gradient expansion. The fact that a gap does not appear implies that the NG modes are massless and that dissipation is always higher order in the derivative expansion. Thus, the effective action of the hydrodynamical NG modes is subject to extra constraints compared to that of a general QFT.
@@
$P\to 0$ limit corresponds to the nonrelativistic regime of the UV sector. In this case, we find $\omega_{0}=0$, and then
$\omega_{1}=-\frac{ik^{2}(3\rho\lambda_{2}+(3\zeta+4\eta)\rho_{\beta})}{3\rho\beta\rho_{\beta}},$ (5.57)
@@
which can be also found when taking $P\to 0$ and $P_{\beta}\to 0$ in Eq. (5.54). Thus, for a vanishing pressure limit, the NG modes are purely diffusive. $\omega_{1}$ depends on $\lambda_{2}$ and $\zeta$ in this limit, which are now hydrodynamical invariants as $\tilde{\lambda}\to\lambda_{2}$ and $\tilde{\zeta}\to\zeta$ in $P\to 0$ limit. $\lambda_{1}$ and $\epsilon$ are eliminated by field redefinition, and the nontrivial first-order correction in the energy-momentum tensor is only spatial.
@@ 5.4 Longitudinal mode for a general charged fluid
For a general charged fluid, the mode analysis is more complicated than the neutral case. The dispersion relation is now 6th order algebraic equation (for a conformal U(1) charged fluid, see Ref. [31])
$0=\sum_{n=0}^{6}d_{n}\omega^{n},$ (5.58)
@@
$\left.+(\beta P_{\beta}-\mu P_{\mu})\left(3\beta n\tau_{3}-3\beta^{2}\chi_{1}(\rho+P)\right)\right),$ (5.59)
@@
$\left.-i(3\zeta+4\eta)\left(\beta\left(\beta^{2}(-\rho_{\beta})\chi_{1}+\beta\mu\rho_{\mu}\chi_{1}+\beta n_{\beta}\tau_{3}-\rho_{\mu}\tau_{1}\right)+n_{\mu}(\kappa_{1}-\beta\mu\tau_{3})\right)\right)$ (5.60)
@@
$+3(\rho+P)(\beta(\beta\chi_{1}(\beta\rho_{\beta}-\mu\rho_{\mu})+\beta\nu_{5}n_{\beta}+\rho_{\mu}\tau_{1})+n_{\mu}(\lambda_{2}-\beta\mu\nu_{5})))$ (5.61)
@@
$+\beta\chi_{1}(\beta\kappa_{2}\rho_{\beta}-\kappa_{2}\mu\rho_{\mu}+\rho(-\epsilon))+\beta\mathcal{P}(\lambda_{2}\nu_{3}-\nu_{1}\nu_{5}+\nu_{4}\tau_{1}-\chi_{1}\epsilon))))$ (5.62)
@@
$+\kappa_{1}\nu_{4}(\nu_{2}+\tau_{2})+\lambda_{1}\nu_{1}\tau_{3}+\kappa_{2}(-\lambda_{2}\nu_{3}+\nu_{1}\nu_{5}-\nu_{4}\tau_{1}+\chi_{1}\epsilon)-\nu_{2}\tau_{3}\epsilon-\tau_{2}\tau_{3}\epsilon))$ (5.63)
@@
$+\beta(\beta\nu_{3}(\mu\rho_{\mu}-\beta\rho_{\beta})-\nu_{1}\rho_{\mu}+\beta\nu_{4}n_{\beta}))$ (5.64)
$d_{6}$ $=3\beta^{2}\kappa_{2}(\nu_{3}\epsilon-\nu_{1}\nu_{4})$ (5.65)
As in the case of the neutral fluid, the above dispersion relation is reliable up to $\mathcal{O}(\varepsilon)$, and we need to solve the dispersion relation order by order in $\varepsilon$. In the low energy limit, the zeroth order solution is found as
$\omega_{0}=\pm c_{s/n}k,~c_{s/n}^{2}\equiv\left(\frac{\partial P}{\partial\rho}\right)_{s/n}.$ (5.66)
The adiabatic sound velocity $c_{s/n}$ contains Eq. (5.53) in the neutral limit. The explicit form of $c_{s/n}$ is found in Eq. (B.13). Zeroth order thermodynamics is stable if $c_{s/n}^{2}\geq 0$. The first-order solution is found as
$\omega_{1}$ $=-i\Gamma_{\rm L}k^{2},$ (5.67)
$\Gamma_{\rm L}$ $\equiv\frac{1}{2(\rho+P)}\left[\frac{4}{3}\eta+{\bm{\zeta}}+f^{2}\left(\frac{\partial P}{\partial\rho}\right)_{s/n}{\bm{\chi}}_{1}+\beta\left(\frac{\partial P}{\partial n}\right)_{\rho}{\bm{\tau}}\right],$ (5.68)
where ${\bm{\zeta}}$, ${\bm{\chi}}_{1}$ and ${\bm{\tau}}$ are defined in Sec. 3, and we defined
$f^{2}$ $=\frac{\beta^{2}(\rho+P)^{2}((\rho+P-\mu n)(\beta n_{\beta}-\mu n_{\mu})-\beta n\rho_{\beta})^{2}}{\left(n_{\mu}(\rho+P-\mu n)^{2}-\beta n(n_{\beta}(\mu n-2(\rho+P))+n\rho_{\beta})\right)^{2}}.$ (5.69)
This is the main result of this paper. Although Eq. (5.68) is written in a simple form, concrete expressions for $\left(\frac{\partial P}{\partial\rho}\right)_{s/n}$ and $\left(\frac{\partial P}{\partial n}\right)_{\rho}$ are tedious, as shown in App. B.3. Note that $f^{2}\geq 0$ since the thermodynamic variables are real. ${\bm{\chi}}_{2}$ drops from the first order solution. As the bold faces are hydrodynamical frame invariants, the mode analysis does not depend on the hydrodynamical frame choice. We have substituted $\omega_{0}$ into the first-order dispersion relation to find $\omega_{1}$, which corresponds to solving the on-shell hydrodynamics for NG modes. This is why the boldfaces write the first-order dispersion relation. Importantly, there is no ambiguity in the use of the on-shell condition, in contrast to the case in Sec. 3. As discussed in Sec. 3, the KMS condition and unitarity of the UV system implies ${\bm{\chi}}\geq 0$, ${\bm{\tau}}=0$ and ${\bm{\zeta}}\geq 0$. Therefore, $\Gamma_{\rm L}\geq 0$ and stable if $\rho+P>0$. Note that we have not taken any specific hydrodynamical frame for writing down the effective action for the NG modes, and hence, this mode analysis applies to the Eckart frame as well.
@@
~$\tilde{\mathcal{E}}^{\rm L}=k^{3}\left(\frac{4}{3}\eta+\zeta+\lambda_{1}+\frac{n(\lambda_{2}+\epsilon)}{\beta n_{\beta}-\mu n_{\mu}}\right)+\mathcal{O}(\varepsilon^{2}).$ (5.70)